# Variational Source-Channel Coding for Semantic Communication

### v2

@@ II-C Variational Source-Channel Coding for Channel Matching
By integrating AI principles with communication theory, the VSCC method establishes a robust algorithmic foundation for training semantic communication systems. As shown in Fig. 3, the VSCC method aims to recover the original source data distribution, which is denoted as $p_{x}(x)$. The received data distribution is represented as $q_{x}(x)$. This objective is achieved through two steps.
@@
$\geq$ $\text{KL}\left(p_{x}(x)\|q_{x}(x)\right),$
@@
$\quad+\mathbb{E}_{z\sim p_{z\mid x}(z\mid x)}\left[-\log q_{x\mid z}(x\mid z)\right]\big{\}},$
@@
It follows that $p_{z|x}(z|x)=\mathcal{N}(\mu=\mu_{1},\sigma^{2}=\sigma_{1}^{2}+\sigma_{2}^{2})$. Additionally, we assume that the true latent variable $z$ follows a distribution similar to the channel noise. Specifically, let $q_{z}(z)=\mathcal{N}(0,\sigma_{2}^{2}+d)$, where $d$ is a channel matching constant (CMC). The CMC is assumed to represent the true variance of the encoded vector $y$. In order to generate source coding vectors according to channel SNR, CMC should be adjusted based on the SNR of the channel. Substituting $p_{z|x}(z|x)$ and $q_{z}(z)$ into Eq. 7 yields:
@@
Thus, we can get the loss function of the VSCC method:
@@
On one hand, unlike VAE, the VSCC method uses the first term in Eq. 10 to align the latent variable $z$ as closely as possible with the channel noise distribution, rather than with a standard Gaussian distribution. This alignment implies that the latent variables $z$ are derived from the original data $x$ by superimposing channel noise. As a result, the channel is no longer an obstacle to be overcome but becomes an integral part of the encoding process. The channel noise aids in eliminating redundancy in $x$, leading to different feature distributions represented by the latent variables $z$. If these feature distributions are considered to be the semantic information of the original message distribution, then the VSCC method effectively achieves both the elimination of data redundancy and the extraction of semantic information.
On the other hand, unlike AE, the VSCC method does not require the original data $x$ to be identical to the received data $\hat{x}$. Instead, it utilizes the second term of Eq. 10 to ensure that the original data distribution and the received data distribution are as similar as possible. This approach not only constrains the encoder to encode based on the distributional features of the original message, resulting in specific feature distributions, but also guides the decoder to reconstruct data that aligns with the distributional features of the original message. This dual consideration ensures that both the encoder and decoder are optimized for semantic fidelity, within the constraints imposed by the channel’s nature.
@@ III Architecture and Implementations
III-A Semantic Communication Model
@@
The image to be transmitted, denoted as $X\subset\mathbb{R}^{H\times W\times C}$, has pixel values ranging from 0 to 255. Since DNNs typically process data within the [0, 1] range to align different data features and accelerate convergence, $X=\{x_{i};i=1,2,\dots,H\times W\times C\}$ is first normalized. The normalized image then enters the joint encoder $f_{\text{en}}(x;\theta)$, where $\theta$ represents the trainable parameters. The joint encoder initially uses a Convolutional Neural Network (CNN) layer to increase the feature dimension $C$ of the image to 32. Subsequently, the image passes through four ResNet blocks and DownSample layers. In each cycle, the ResNet block increases the feature dimension $C$ to 32, 64, 128, and 192, while the DownSample layers halve the height $H$ and width $W$ of the image. The DownSample layers are implemented using CNNs, compressing the image size and extracting its features. Each cycle processes images with different heights $H$, widths $W$, and feature dimensions $C$, enabling feature extraction across various dimensions. Following this, three additional ResNet blocks and an Attention block capture the remaining feature information in the 192-dimensional space. Finally, a CNN layer compresses the normalized feature information into a encoded vector $Y$, which can be regarded as following the distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$. This encoded vector are then transmitted through the channel.
@@ IV Results
IV-A Semantic communication capability of VSCC method
@@
IV-B Channel matching capability of VSCC method
@@
IV-C Comparison of VSCC, VAE, and AE Methods
@@
V Conclusion and Future Work
@@
This paper identifies three key challenges in semantic communication. To address these issues, we analyze the differences between lossless and lossy communication, highlighting that for lossy communication, the assumptions required for the joint source-channel coding theorem are broken, rendering the separation theorem invalid. In this context, JSCC emerges as the optimal transmission strategy. Further, we highlight that the primary distinction between semantic communication and classical communication lies in whether the data is lossy. On the one hand, based on research in semantic information, it is clear that while semantics inherently depend on the existence of data, not all data necessarily contribute to semantics. Consequently, the goal of semantic encoding is to eliminate redundant data and extract its semantic features. The semantic communication significantly reduces the amount of data that needs to be transmitted, which is a major source of its gain. On the other hand, the presence of a knowledge base and specific communication tasks in semantic communication further expands beyond the boundary conditions of CIT, allowing joint coding to achieve optimal transmission while further reducing the transmission data, which constitutes another key advantage of semantic communication.
@@
As semantic communication requires a joint coding method that permits data distortion, we propose an innovative VSCC method. By introducing latent variables, the channel is integrated into the joint coding process. VSCC encodes original data into a channel-dependent feature distribution, allowing data to be categorized by the encoded feature distributions. This method enables data distortion while preserving the original data’s distributional characteristics. Unlike AE methods, VSCC focuses on lossless transmission of the original data distribution rather than the data itself. We constructed the JSCC model using ResNet Blocks and trained it with VSCC, VAE, and AE methods. Experiments show that VSCC model effectively extracts common features from images, outputting them as the variances of the feature distributions. These variances, representing data dispersion, align with lossy transmission. Further, it can be found that the variances of feature distributions are influenced by the training channel SNRs, and varying CMCs during training leads to different transmission performances. It demonstrates that the channel is integrated into the production of feature distributions in VSCC, becoming part of the joint encoder rather than an obstacle. The VSCC model outperforms the VAE model in data recovery but falls short of the AE model, which targets lossless transmission. While an objective semantic evaluation metric is lacking, the comparison of PSNR and SSIM suggests that VSCC method achieves a degree of semantic recovery, with transmitted data more closely aligning with human perception.

### v3

@@ III-C Variational Source-Channel Coding for Channel Matching
By integrating variational inference and VIB principle with communication theory, the VSCC method establishes a robust algorithmic foundation for training semantic communication systems. As shown in Fig. 3, the VSCC method aims to recover the original source data distribution, which is denoted as $p_{x}(x)$. The received data distribution is represented as $q_{x}(x)$. This objective is achieved through two steps.
@@
$\geq$ $\text{KL}\left(p_{x}(x)\|q_{x}(x)\right).$
@@
$\quad+\mathbb{E}_{z\sim p_{z\mid x}(z\mid x)}\left[-\log q_{x\mid z}(x\mid z)\right]\big{\}}.$
It is worth noting that the Eq. 9 is nearly identical to the VIB Eq. 4, with the only difference being the weighting factor $\beta$. This confirms the equivalence between variational inference and the VIB principle. The weight $\beta$ controls the degree of information compression in the VIB, which will be reflected in the second step through the CMC.
@@
It follows that $p_{z|x}(z|x)=\mathcal{N}(\mu=\mu_{1},\sigma^{2}=\sigma_{1}^{2}+\sigma_{2}^{2})$. Additionally, we assume that the true latent variable $z$ follows a distribution similar to the channel noise. Specifically, let $q_{z}(z)=\mathcal{N}(0,\sigma_{2}^{2}+d)$, where $d$ is the CMC. The CMC is assumed to represent the true variance of the encoded vector $y$. At the same time, the CMC also reflects the role of the parameter $\beta$ in Eq. 4 of the VIB in controlling the degree of information compression, with the distinction that $\beta$ operates multiplicatively, whereas the CMC functions additively.
In order to generate source coding vectors according to channel SNR, CMC should be adjusted based on the SNR of the channel. Substituting $p_{z|x}(z|x)$ and $q_{z}(z)$ into Eq. 9 yields:
@@
Thus, we can get the loss function of the VSCC:
@@
On one hand, unlike VAE, the VSCC method uses the first term in Eq. 12 to align the latent variable $z$ as closely as possible with the channel noise distribution, rather than a standard Gaussian distribution. This alignment implies that the latent variables $z$ are derived from the original data $x$ by superimposing channel noise. As a result, the channel is no longer an obstacle to be overcome but becomes an integral part of the encoding process. The channel noise aids in eliminating redundancy in $x$, leading to different feature distributions represented by the latent variables $z$. If these feature distributions are considered to be the semantic information of the original message distribution, then the VSCC method effectively achieves both the elimination of data redundancy and the extraction of semantic information.
On the other hand, unlike AE, the VSCC method does not require the original data $x$ to be identical to the received data $\hat{x}$. Instead, it utilizes the second term of Eq. 12 to ensure that the original data distribution and the received data distribution are as similar as possible. This approach not only constrains the encoder to encode based on the distributional features of the original message, resulting in specific feature distributions, but also guides the decoder to reconstruct data that aligns with the distributional features of the original message. This dual consideration ensures that both the encoder and decoder are optimized for semantic fidelity, within the constraints imposed by the channel.
@@ IV Architecture and Implementations
IV-A Semantic Communication Model
@@
The image to be transmitted, denoted as $X\subset\mathbb{R}^{H\times W\times C}$, has pixel values ranging from 0 to 255. Since DNNs typically process data within the [0, 1] range to align different data features and accelerate convergence, $X=\{x_{i};i=1,2,\dots,H\times W\times C\}$ is first normalized. The normalized image then enters the joint encoder $f_{\text{en}}(x;\theta)$, where $\theta$ represents the trainable parameters. The joint encoder initially uses a CNN layer to increase the feature dimension $C$ of the image to 32. Subsequently, the image passes through four ResNet blocks and DownSample layers. In each cycle, the ResNet block increases the feature dimension $C$ to 32, 64, 128, and 192, while the DownSample layers halve the height $H$ and width $W$ of the image. The DownSample layers are implemented using CNNs, compressing the image size and extracting its features. Each cycle processes images with different heights $H$, widths $W$, and feature dimensions $C$, enabling feature extraction across various dimensions. Following this, three additional ResNet blocks and an Attention block capture the remaining feature information in the 192-dimensional space. Finally, a CNN layer compresses the normalized feature information into a encoded vector $Y$, which can be regarded as following the distribution $\mathcal{N}(\mu_{1},\sigma_{1}^{2})$. This encoded vector are then transmitted through the channel.
@@ V Results
V-A Semantic communication capability of VSCC method
@@
V-B Channel matching capability of VSCC method
@@
V-C Comparison of VSCC, VAE, and AE Methods
@@
VI Conclusion and Future Work
VI-A Conclusion
@@
This paper identifies three key challenges in semantic communication. To address the first problem, we highlight that the primary distinction between semantic communication and classical communication lies in whether the data is lossy. Based on research in semantic information, it is clear that while semantics inherently depend on the existence of data, not all data necessarily contribute to semantics. Consequently, the goal of semantic encoding is to eliminate redundant data and extract its semantic features. The semantic communication significantly reduces the amount of data that needs to be transmitted, which is a major source of its gain.
For the second problem, semantic communication extends beyond the boundary conditions of CIT by incorporating a knowledge base and specific communication tasks. This enables JSCC to achieve optimal transmission while further reducing data volume. Building upon the IB principle, we can establish a semantic communication system model within the JSCC framework. Moreover, by implementing VIB, we can integrate deep learning into the practical training process of semantic communication systems. However, the VIB characterizes the process of information compression and task completion from an information-theoretic perspective, yet it does not specify concrete implementation methods for encoding and decoding. Therefore, we introduce variational inference to conduct detailed modeling of the encoder and decoder that incorporate the channel.
At last, we propose an innovative VSCC approach by incorporating channel modeling into the encoder, which simultaneously provides the solution to the third problem. Unlike VAE, which deterministically presumes a predefined distribution for latent variables, VSCC dynamically derives the latent distribution by integrating channel modeling into the encoder. This approach naturally induces channel-adaptive latent distributions, with theoretical analysis revealing their dynamic variation with channel conditions. Furthermore, as a communication-oriented framework, VSCC introduces CMC as a tunable hyperparameter in variational inference to balance channel noise against source distortion. Future work will focus on developing a learnable component within the VIB framework to adaptively control information compression, thereby achieving better channel adaptation. In contrast, VAE remains solely a generative model without such transmission optimization mechanisms.
@@
We constructed the semantic communication system using ResNet Blocks and trained it with three approaches: VSCC, VAE, and AE methods. Experiments show that the VSCC model effectively extracts common features from images, outputting them as the variances of the feature distributions. These variances, representing data distortion, align with lossy transmission. Further, it can be found that the variances of feature distributions are influenced by the training channel SNRs, and varying the CMC during training leads to different transmission performances. It demonstrates that the channel is integrated into the production of feature distributions in the VSCC, becoming part of the joint encoder rather than an obstacle. The VSCC model outperforms the VAE model in data recovery but falls short of the AE model, which targets lossless transmission. While an objective semantic evaluation metric is lacking, the comparison of PSNR and SSIM suggests that the VSCC method achieves a degree of semantic recovery, with transmitted data more closely aligning with human perception.
VI-B Future Work